# TripleMixer: A 3D Point Cloud Denoising Model for Adverse Weather

### v1

@@ V-D Quantitative Result
Results On Weather-NuScenes: We also compared the denoising performance of our model with six state-of-the-art methods across three scenarios in the Weather-NuScenes dataset. As shown in Table LABEL:Weather-NuScenes, our TripleMixer model achieved the highest average mIoU of 97.21 across these three scenarios. Notably, the average mIoU of our model is 6.9% higher than 3D-OutDet[22] and 15.20% higher than 4DenoiseNet[21]. Additionally, our model achieved the best mIoU accuracy in each scenario (Snow, Fog, and Rain) compared to other denoising models, with mIoUs of 95.78, 97.42, and 98.48, respectively. These results demonstrate that our model exhibits superior performance in the Weather-NuScenes dataset across three different adverse weather scenarios compared to other denoising models.
We further present the mIoU performance of our model and three other denoising models across six scenarios in the Weather-KITTI and Weather-NuScenes datasets, as shown in Figure 7. Our model not only achieved the highest mIoU but also exhibited the smallest variance in mIoU accuracy across the six scenarios, with a variance of only 0.96. This indicates that our denoising model has strong generalization capability and is the most robust compared to the other denoising models, effectively denoising across diverse weather conditions.
Visualization: We visualized our model’s prediction results for point cloud denoising performance on WADS[17], Weather-KITTI, and Weather-NuScenes, compared with those of the state-of-the-art models, including: 4DenoiseNet[21], 3D-OutDet[22], and Cylinder3D[33]. As illustrated in Figure 8, it is evident that our TripleMixer model delivers more accurate predictions, achieving optimal denoising results across different weather scenarios compared to the other three denoising models.
VI Discussions
VI-A Ablation Study:
In this section, we perform thorough ablation studies on the point cloud denoising task to investigate the effect of each components, projection plane resolution, wavelet decomposition levels and loss functions of our TripleMixer model. All ablation experiments are conducted on the real-world WADS[17] dataset, with same experimental settings as in Section V.
Table VII: Ablation Results of Each Component. GMX: Geometry Mixer Layer, FMX: Frequency Mixer Layer, and CMX: Channel Mixer Layer.
@@ VI-A Ablation Study:
Table VIII: Ablation Results of Different XY Directions Projection Resolution When Z Direction Projection Resolution is set to 32
@@
Table IX: Ablation Results of Different Z Direction Projection Resolution When XY Directions Projection Resolution is set to [256,256]
@@
Table X: Ablation Results of Different Wavelet Decomposition Levels in the Frequency Mixer Layer
@@
Table XII: Semantic Segmentation Performance Comparison on WADS Datasets. The Row Marked with $\Delta$ Represents the Performance Difference of Our Denoising Model TripleMixer for Downstream Semantic Segmentation Tasks.
Methods mIoU
car
truck
other-vehicle
person
road
parking
sidewalk
other-ground
building
fence
other-structure
vegetation
trunk
terrain
pole
traffic-sign
other-objec
acc-snow
SalsaNext[32] 14.32 45.98 4.84 0.00 0.00 50.56 4.21 6.53 2.47 46.85 2.40 1.05 47.35 0.34 0.31 6.12 7.46 0.29 29.91
SalsaNext[32] + TripleMixer 15.61 50.22 1.20 0.45 0.93 53.83 14.67 8.79 7.92 54.97 3.86 0.50 43.25 0.29 0.35 7.08 11.00 1.41 20.22
$\Delta$ +1.29 +4.24 -3.64 +0.45 +0.93 +3.27 +10.46 +2.26 +5.45 +8.12 +1.46 -0.55 -4.10 -0.05 +0.04 +0.96 +3.54 +1.12 -9.69
Cylinder3D[33] 19.40 61.88 0.00 0.00 0.00 47.45 0.15 4.46 0.17 72.01 16.03 1.01 51.76 0.00 0.00 24.40 14.42 0.36 54.16
Cylinder3D[33] + TripleMixer 21.03 66.03 0.00 0.26 0.48 46.90 2.20 5.99 0.57 79.13 14.47 1.37 57.80 0.18 0.24 31.19 17.92 1.62 52.28
$\Delta$ +1.63 +4.15 0.00 +0.26 +0.48 -0.55 +1.05 +1.53 +0.40 +7.12 -1.56 +0.36 +6.04 +0.18 +0.24 +6.79 +3.50 +1.26 -1.88
PolarNet[50] 19.28 59.10 0.03 0.00 0.00 48.25 0.01 0.14 0.20 72.30 3.84 0.01 61.94 0.00 0.00 19.94 21.77 0.39 59.27
PolarNet[50] + TripleMixer 20.36 60.21 0.08 0.12 0.06 59.38 0.01 0.01 0.04 68.40 13.29 0.42 59.08 0.22 0.26 27.45 21.78 2.32 53.23
$\Delta$ +1.08 +1.11 +0.05 +0.12 +0.06 +11.13 0.00 -0.13 -0.16 -3.90 +9.45 +0.41 -2.86 +0.22 +0.26 +7.51 +0.01 +1.93 -6.04
@@
Effects of Each Component. To testify the effectiveness of each component of TripleMixer, we conduct an extensive ablation study and list the result in Table VII. As shown in Table VII, Experiments I-III represent the accuracy of our TripleMixer model after sequentially removing the GMX Layer, the FMX Layer, and the CMX Layer, respectively. From Experiments I and IV, we observe that the GMX Layer contributes significantly, resulting in F1 and mIoU scores of 0.49 and 0.90, respectively. Similarly, Experiments II and IV indicate that the FMX Layer enhances performance, yielding F1 and mIoU improvements of 0.65 and 1.19, respectively. Finally, the impact of the CMX Layer is evident from Experiments III and IV, showing F1 and mIoU increases of 0.12 and 0.25, respectively. These results demonstrate that the GMX Layer, the FMX Layer, and the CMX Layer are all crucial for improving model performance, each contributing significantly to the overall effectiveness of our TripleMixer model.
Effects of Projection Plane Resolution. In the Frequency Mixer Layer, the 3D geometric features are projected onto three 2D planes along the X, Y, and Z axes. An adaptive wavelet decomposition is then performed on these projected 2D planes to extract multi-scale features. The initial projection resolution of each 2D plane influences the effectiveness of the multi-scale feature extraction. In this part, we conduct ablation experiments to assess the impact of different projection plane resolutions on the performance of our TripleMixer model. Due to the different horizontal and vertical emission ranges of LiDAR, we set different projection resolutions for the XY directions and the Z direction. We first set the projection resolution for the Z direction to 32 in order to select the optimal projection resolution for the XY directions. Table LABEL:plane_reso_xy shows the impact of different projection resolutions in the XY directions on the model’s denoising performance, with the Z direction projection resolution set to 32. As shown in Table LABEL:plane_reso_xy, the optimal mIoU accuracy is achieved when the projection resolution in the XY directions is set to [256,256]. Both excessively high resolutions such as [512,512] and excessively low resolutions such as [64,64] result in suboptimal performance. Therefore, we selected the resolution of [256,256] for the XY directions. We further verified the impact of different resolutions in the Z direction on model accuracy. The results are shown in Table LABEL:plane_reso_z, where we experimented with five different resolutions for the Z direction. As indicated in Table LABEL:plane_reso_z, with the XY directions’ resolution fixed, the optimal mIoU accuracy is achieved when the Z direction resolution is set to 32. Similar to the results for the XY directions, excessively high or low resolutions in the Z direction also resulted in poorer accuracy. Therefore, in our practical experiments, we chose the projection resolution of [256,256,32] for the XYZ directions.
Effects of Wavelet Decomposition Levels. In this part, we investigate the impact of varying levels of wavelet decomposition in the Frequency Mixer Layer on our model’s performance. With the projection resolution set to [256,256,32] for the XYZ directions, we select a maximum wavelet decomposition level of 3. This choice is due to the fact that if the wavelet decomposition level exceeds 3, the resolution in the Z direction becomes $\leq 2$, which hinders the extraction of effective information. Consequently, we analyze the model’s performance across a range of wavelet decomposition levels, from none to the maximum level of 3. As shown in Table LABEL:wavelet, not employing any wavelet decomposition levels results in the poorest model accuracy, highlighting the critical role that wavelet transformations play in optimizing model performance. Additionally, utilizing the maximum number of decomposition levels also leads to suboptimal performance, as excessive wavelet decomposition fails to extract more effective information. Our TripleMixer achieves the best overall accuracy when the decomposition level is set to 2.
@@
Effects of Loss Functions. Our model’s loss function primarily consists of three components: the cross-entropy loss, the Lovász loss, and the wavelet regularization terms. We use the cross-entropy loss as the baseline to test the impact of different loss functions on model accuracy. As shown in Table XI, after adding only the Lovász loss, the mIoU and F1 scores of our TripleMixer model improved by 0.37 and 0.21, respectively. Additionally, after incorporating only the wavelet regularization terms, the model’s mIoU and F1 scores increased by 0.26 and 0.14, respectively. When using the combination of all three loss functions, our denoising model achieved the highest mIoU and F1 scores.
VI-B Experiments on Segmentation Tasks
In this section, we apply the proposed denoising model to LiDAR-based downstream perception tasks. We conducted point cloud semantic segmentation experiments on the WADS dataset, both with and without our denoising model. These experimental results demonstrate the generalization ability of our proposed method and its effectiveness on LiDAR point cloud processing.
The performance of the point cloud semantic segmentation task for SalsaNext[32], Cylinder3D[33] and PolarNet[50] models before and after removing snow with our TripleMixer model is shown in Table LABEL:wads_denoise, where $\Delta$ indicates the performance differences resulting from the snow removal. As illustrated in Table LABEL:wads_denoise, with the assistance of our model, the semantic segmentation performance of the SalsaNext[32], Cylinder3D[33], and PolarNet[50] models has significantly improved. The performance enhancements are quantified as increases of 1.29%, 1.63%, and 1.08%, respectively. Notably, before the snow removal by TripleMixer, none of the three models were capable of detecting the ’other-vehicle’ and ’person’ categories. However, with the integration of our model, there is a marked improvement in the detection accuracy for these categories. This demonstrates that our model is beneficial and versatile for LiDAR-based downstream perception tasks.
VII Conclusion and Future Work
In this work, we first introduce two large-scale adverse weather point cloud datasets, Weather-KITTI and Weather-NuScenes, featuring a wide range of weather scenarios, complete original LiDAR information, and point-wise weather and semantic labels. These datasets provide robust data support for further research into adverse weather conditions. Subsequently, we propose a novel point cloud denoising model, the TripleMixer network, which integrates spatial geometric features, frequency domain features, and multi-channel contextual information to achieve high denoising accuracy. This model demonstrates optimal performance on both real-world datasets and our newly proposed datasets. Finally, we validate the significant performance enhancement of our denoising model for downstream point cloud tasks, specifically semantic segmentation, under adverse weather conditions. In future work, we aim to incorporate additional sensors, such as cameras and IMUs, to further explore the perception capabilities of multi-sensor fusion in extreme weather conditions.

### v2

@@ V-A Denoising Benchmark
Results On Weather-NuScenes: We also compared the denoising performance of our model with seven state-of-the-art methods across three scenarios in the Weather-NuScenes dataset. As shown in Table LABEL:Weather-NuScenes, our TripleMixer model achieved the highest average mIoU of 97.21 across these three scenarios. Notably, the average mIoU of our model is 6.9% higher than 3D-OutDet[14] and 15.20% higher than 4DenoiseNet[21]. Additionally, our model achieved the best mIoU accuracy in each scenario (Snow, Fog, and Rain) compared to other denoising models, with mIoUs of 95.78, 97.42, and 98.48, respectively. These results demonstrate that our model exhibits superior performance in the Weather-NuScenes dataset across three different adverse weather scenarios compared to other denoising models.
Figure 8 visualizes denoising results on WADS [18], Weather‑KITTI, and Weather‑NuScenes, showing that TripleMixer consistently produces more accurate and cleaner reconstructions than 3D‑OutDet [14] and LSK3DNet [46] across various weather conditions.
Table VIII: Benchmarking Results for the SS Task (%). All SS models are pretrained on SemanticKITTI [36] and inferred on SemanticSTF [23]. Denoising models are trained on Weather-KITTI and used as preprocessing during inference.
Methods mIoU car bi.cle mt.cle truck oth-v. pers. bi.list mt.list road parki. sidew. oth-g. build. fence veget. trunk terra. pole. traf.
$S$ Preprocessing
SphereFormer [47] Baseline 11.26 19.12 0.00 0.00 13.83 4.12 0.99 6.75 0.00 22.28 4.69 15.80 4.43 25.53 13.27 24.23 4.38 37.74 14.71 1.99
4DenoiseNet [21] 11.03 18.24 0.00 0.00 13.42 3.92 1.00 2.17 0.00 22.01 4.72 16.25 4.74 25.33 14.68 23.97 4.50 38.15 14.57 1.98
3D-OutDet [14] 11.87 23.26 0.00 0.00 15.94 3.58 1.64 1.54 0.00 24.44 4.68 17.41 5.73 25.04 13.99 27.25 5.30 38.57 15.20 2.04
TripleMixer (Ours) 13.71 27.19 0.00 0.00 21.20 4.07 0.21 5.68 0.00 26.72 4.76 19.68 8.85 23.97 33.64 22.86 4.31 40.95 14.48 1.84
SFPNet [48] Baseline 11.14 14.36 0.00 0.00 11.60 0.00 0.23 5.29 0.00 38.06 4.33 13.71 5.09 17.19 12.06 30.01 4.97 37.98 14.86 1.92
4DenoiseNet [21] 11.07 14.01 0.00 0.00 10.89 0.00 0.24 4.55 0.00 38.18 4.12 13.73 5.08 17.26 12.39 29.73 4.96 38.07 15.10 1.95
3D-OutDet [14] 12.12 23.51 0.00 0.00 11.27 0.00 0.34 4.14 0.00 43.87 6.45 14.83 5.81 17.30 11.76 30.38 4.88 38.37 15.42 2.01
TripleMixer (Ours) 13.23 21.33 0.06 0.00 15.95 0.00 0.36 4.91 0.00 42.35 8.37 15.74 6.58 19.88 18.61 37.01 5.05 38.00 15.39 1.87
PTv3 [49] Baseline 14.58 15.27 1.13 0.00 3.56 0.40 0.21 0.00 37.25 44.10 12.05 9.95 6.46 25.30 25.70 37.46 8.63 36.97 10.10 2.46
4DenoiseNet [21] 14.53 14.88 1.12 0.00 3.60 0.48 0.21 0.00 37.25 44.07 11.86 9.92 6.43 25.42 25.95 37.57 8.70 37.00 9.22 2.48
3D-OutDet [14] 16.40 19.87 2.14 0.00 9.58 2.48 0.21 0.00 40.56 44.51 11.30 9.87 8.48 25.28 30.03 40.60 8.54 42.18 12.12 2.79
TripleMixer (Ours) 17.18 23.64 2.80 0.00 19.88 8.26 1.74 0.00 42.87 44.06 11.36 9.91 9.36 29.09 28.53 36.93 9.26 36.93 9.87 1.98
Figure 9: Qualitative results of the SS benchmark under adverse weather. The “Denoising” row presents outputs of different denoising models, while the “Semantic Segmentation” row shows PTv3[49] results on the corresponding denoised point clouds. Red points denote weather‑induced noise, and blue points represent the “car” class.
V-B Semantic Segmentation (SS) Benchmark
Benchmark Setting: We evaluate the downstream SS models on the real‑world adverse‑weather SemanticSTF [23] dataset, which provides semantic annotations for 19 categories consistent with SemanticKITTI. To comprehensively assess performance, we select three recent state‑of‑the‑art semantic segmentation models: SphereFormer [47], SFPNet [48], and Point Transformer v3 (PTv3) [49]. In our experiments, they pretrained on SemanticKITTI under clear‑weather conditions are directly evaluated on SemanticSTF. Furthermore, to examine the impact of denoising on SS models performance, we incorporate three specialized denoising networks: 4DenoiseNet[21], 3D‑OutDet [14], and our proposed TripleMixer, all trained in a supervised manner on the our Weather‑KITTI datasets and directly generalized to SemanticSTF. Evaluation follows standard protocols, computing the Intersection‑over‑Union (IoU) for each class and reporting the mean IoU (mIoU) across all classes.
Benchmark Results: Table LABEL:LSS_BENCH presents the benchmarking results for the SS task under real-world adverse weather on the validation set of SemanticSTF [23] dataset. The baseline denotes direct application of SS models without denoising preprocessing. Across all segmentation backbones, including SphereFormer [47], SFPNet [48], and PTv3 [49], the proposed TripleMixer consistently yields the highest mIoU, demonstrating significant improvements over both the baseline and other denoising methods. Specifically, when integrated with SphereFormer, TripleMixer achieves an mIoU of 13.71, improving by 2.45 points over the baseline. Similar gains are observed with SFPNet and PTv3, where TripleMixer improves the mIoU by 2.09 and 2.60 points, respectively. On average, TripleMixer achieves a 2.38-point improvement across all three backbones, significantly surpassing the performance of 3D-OutDet, which only provides an average gain of 1.13 points. Notably, 4DenoiseNet, which demonstrates limited denoising effectiveness on our proposed datasets, fails to improve real-world segmentation performance and even leads to performance degradation compared to the baseline. At the category level, TripleMixer delivers substantial gains in key semantic classes, improving IoU by 7.81 for car and 9.35 for truck over the baseline, averaged across the three segmentation models. It also enhances performance in challenging structural categories such as buildings and fences. These results underscore the robustness and practical utility of TripleMixer and our dataset for improving LiDAR semantic segmentation in real‑world conditions. We visualize the SS benchmark under adverse weather in Figure 9 to illustrate denoising results in real‑world scenarios and the subsequent semantic segmentation performance, focusing on the key semantic class “car”. Compared with 3D‑OutDet, TripleMixer achieves more effective denoising and enhances segmentation performance under real‑world extreme conditions.
Table IX: Benchmarking Results for the PR Task. All PR models are trained on the KITTI [35] and inferred on Boreas (BR) [41] and Weather-KITTI (WK). Denoising models, trained on Weather-KITTI, serve as preprocessing during inference.
Method mRS WK Snow WK Fog WK Rain BR Easy BR Hard
PR Preprocessing R@1↑ R@1%↑ F1↑ R@1↑ R@1%↑ F1↑ R@1↑ R@1%↑ F1↑ R@1↑ R@1%↑ F1↑ R@1↑ R@1%↑ F1↑
OT [50] Baseline 0.15 0.03 0.26 0.07 0.09 0.33 0.17 0.17 0.52 0.29 0.01 0.11 0.02 0.00 0.09 0.02
4DenoiseNet [21] 0.22 0.03 0.24 0.06 0.21 0.58 0.37 0.25 0.61 0.40 0.01 0.09 0.02 0.00 0.06 0.01
3D-OutDet [14] 0.27 0.07 0.38 0.13 0.22 0.69 0.42 0.27 0.62 0.43 0.02 0.13 0.03 0.01 0.11 0.03
TripleMixer (Ours) 0.31 0.09 0.43 0.17 0.31 0.78 0.47 0.29 0.65 0.44 0.04 0.15 0.06 0.03 0.14 0.05
CVTNet [51] Baseline 0.28 0.03 0.34 0.07 0.29 0.50 0.46 0.19 0.61 0.33 0.14 0.21 0.25 0.09 0.14 0.15
4DenoiseNet [21] 0.33 0.02 0.30 0.03 0.31 0.48 0.46 0.43 0.83 0.60 0.14 0.20 0.25 0.10 0.15 0.17
3D-OutDet [14] 0.37 0.04 0.42 0.09 0.31 0.50 0.47 0.49 0.86 0.65 0.17 0.23 0.26 0.12 0.17 0.19
TripleMixer (Ours) 0.39 0.06 0.44 0.11 0.33 0.54 0.50 0.51 0.89 0.67 0.16 0.29 0.28 0.13 0.24 0.21
LPSNet [52] Baseline 0.30 0.02 0.18 0.04 0.09 0.77 0.16 0.29 0.74 0.46 0.13 0.28 0.26 0.10 0.26 0.19
4DenoiseNet [21] 0.29 0.02 0.22 0.05 0.03 0.77 0.06 0.28 0.72 0.44 0.16 0.28 0.27 0.10 0.24 0.19
3D-OutDet [14] 0.33 0.02 0.27 0.06 0.10 0.85 0.17 0.32 0.75 0.49 0.18 0.29 0.28 0.11 0.27 0.20
TripleMixer (Ours) 0.37 0.03 0.29 0.06 0.14 0.83 0.24 0.37 0.79 0.54 0.15 0.36 0.29 0.11 0.34 0.23
V-C Place Recognition (PR) Benchmark
Benchmark Setting: We selected distinct sequences from the real‑world Boreas [41] dataset and our Weather‑KITTI dataset. In Boreas, the sequence recorded on 2021-11-02-11-16 is used as the database, while the sequence from 2021-01-26-11-22, which features heavy snowfall, serves as the query. Boreas Easy employs only the forward traversal, whereas Boreas Hard includes both directions to introduce greater viewpoint variation. For Weather‑KITTI, we use the snow 00, fog 07, and rain 00 as query sets, with their corresponding clear-weather sequences serving as the database. The benchmark includes three leading place recognition models: OverlapTransformer (OT)[50], CVTNet[51], and LPSNet [52], all pretrained on the KITTI dataset and directly evaluated on Boreas and Weather‑KITTI. The denoising models and training configurations are the same as those used in the SS benchmark.
For the PR evaluation, a match is considered a true recognition if the pose distance is less than 5 meters for the Weather-KITTI, and less than 15 meters for the Boreas dataset. We follow standard metrics widely adopted in prior work [51], including Recall@1 (R@1), Recall@1% (R@1%), and the F1 score. Additionally, we introduce the Mean Robust Score (mRS) to quantify model robustness under adverse weather conditions, defined as:
${\text{mRS}}=\frac{1}{N}\sum\frac{\text{M}_{\text{w}}}{\text{M}_{\text{c}}},\quad\text{M}\in\{\text{R@1},\text{R@1\%},\text{F1}\},$ (17)
where $\text{M}_{\text{w}}$ and $\text{M}_{\text{c}}$ denote the PR evaluation metrics under adverse and clean weather conditions, respectively. $N$ represents the total number of evaluated metrics across all test sequences. A higher value of mRS reflects greater robustness of the PR model when transitioning from clean to adverse weather conditions
Benchmark Results: Table LABEL:LPR_BENCH presents the benchmarking results for the PR task under adverse weather conditions on the Boreas and Weather-KITTI datasets. Across all test sequences, TripleMixer consistently achieves the highest mRS values, demonstrating its effectiveness in enhancing the robustness of PR models under adverse weather conditions. For OT, TripleMixer improves the mRS from the baseline value of 0.15 to 0.31. Similarly, for CVTNet and LPSNet, TripleMixer achieves mRS values of 0.39 and 0.37, respectively, exceeding both the baseline and other denoising models. TripleMixer achieves an average improvement of 56% across three PR models under adverse weather conditions. Notably, it consistently enhances place recognition performance even in the challenging Boreas Hard scenarios. These results highlight the effectiveness of both TripleMixer and the proposed Weather-KITTI dataset in improving the robustness of downstream PR models against real-world adverse weather conditions. Figure 10 visualizes the result of PR benchmark. Most noise points are concentrated near the LiDAR center. Our TripleMixer achieves effective denoising on both the real‑world Boreas dataset and our Weather‑KITTI dataset, thereby improving PR accuracy.
Figure 10: Qualitative results of the PR benchmark under adverse weather. The PR results show queries processed by different denoising methods and their top‑1 retrieved matches using CVTNet [51]. Brown boxes indicate incorrect retrieval, while blue boxes denote correct retrievals.
Table X: Benchmarking Results for the OD Task (%). All OD models are trained on the KITTI dataset [35] and inferred on CADC [29]. Denoising models, trained on Weather-NuScenes, are employed as preprocessing during inference.
Methods mAP Car (3D) Car (BEV) Pedestrian (3D) Pedestrian (BEV)
OD Preprocessing Easy↑ Med↑ Hard↑ Easy↑ Med↑ Hard↑ Easy↑ Med↑ Hard↑ Easy↑ Med↑ Hard↑
TED-S [53] Baseline 12.18 21.85 20.07 15.76 24.16 21.13 18.65 4.26 4.17 3.47 4.54 4.46 3.72
4DenoiseNet [21] 11.91 22.31 20.03 14.88 25.20 21.77 17.51 4.24 3.12 2.81 4.35 3.71 3.01
3D-OutDet [14] 12.63 22.33 20.05 16.66 25.18 22.74 18.50 5.02 4.12 3.64 5.23 4.31 3.80
TripleMixer (Ours) 13.58 24.35 22.06 18.65 27.17 23.73 19.49 5.24 4.20 3.81 5.54 4.73 4.02
PG-RCNN [54] Baseline 12.53 22.36 19.11 16.67 24.91 22.51 19.89 5.36 5.05 1.99 5.41 5.09 2.04
4DenoiseNet [21] 12.25 22.60 19.13 16.72 25.81 22.30 19.73 4.31 4.17 1.59 4.33 4.21 1.84
3D-OutDet [14] 13.43 23.79 19.93 16.11 26.84 23.09 18.96 6.28 6.16 3.80 6.29 6.18 3.84
TripleMixer (Ours) 14.14 24.90 20.23 17.81 28.32 23.43 20.82 6.42 6.39 4.23 6.45 6.43 4.27
VoxT-GNN [55] Baseline 10.02 15.34 14.24 11.93 19.91 18.78 15.38 4.37 4.16 3.71 4.46 4.25 3.79
4DenoiseNet [21] 9.70 15.81 13.84 11.58 20.40 18.23 14.78 3.63 3.48 3.25 4.08 3.70 3.56
3D-OutDet [14] 10.53 16.59 14.94 12.55 21.18 19.31 15.81 4.62 4.18 4.01 4.85 4.26 4.09
TripleMixer (Ours) 12.33 19.41 16.37 14.13 23.99 20.93 18.09 6.41 6.38 4.96 6.14 6.12 4.93
V-D Object Detection (OD) Benchmark
Benchmark Setting: For the OD benchmark, we use the real‑world CADC [29] dataset, which provides object‑level annotations for 3D object detection in a KITTI‑style format, ensuring compatibility with standard evaluation protocols. We evaluate three advanced 3D object detection models, TED‑S [53], PG‑RCNN [54], and VoxT‑GNN [55], using their pretrained versions on the KITTI dataset and directly test them on CADC dataset. Consistent with the SS benchmark, we employ the same denoising models, trained on Weather‑NuScenes and evaluated on CADC, leveraging their shared 32‑beam LiDAR configuration. For evaluation, we follow the KITTI 3D object detection protocol, reporting the mean Average Precision (mAP) over 40 recall positions (R40). We adopt IoU thresholds of 0.7 for cars and 0.5 for pedestrians, while excluding cyclists due to their absence in CADC. Results are reported across three difficulty levels: easy, moderate, and hard.
Benchmark Results: The benchmarking results for the OD task under real-world adverse weather conditions on the validation set of CADC [29] dataset, as summarized in Table LABEL:LOD_BENCH. Evaluations are conducted for both 3D and Bird’s Eye View detection across car and pedestrian categories. The results demonstrate that incorporating TripleMixer as a preprocessing module consistently improves both 3D and BEV detection performance under real-world adverse weather conditions across all OD baselines, including TED-S [53], PG-RCNN [54], and VoxT-GNN [55], with an average performance gain of 16%. Specifically, when applied to PG-RCNN, it increases the overall mAP from 12.53 to 14.14 and improves 3D detection for the Car category under hard conditions from 16.67 to 17.81. Similarly, with VoxT-GNN, it boosts the mAP from 10.02 to 12.33 and enhances Pedestrian 3D detection under hard conditions from 3.71 to 4.96. In contrast, 4DenoiseNet fails to improve detection performance, similar to its results in the SS task. Its limited denoising ability hinders performance on detail-sensitive tasks. These results show that TripleMixer benefits from our Weather-NuScenes dataset, enabling more effective spatial restoration and consistently improving OD performance under real-world extreme weather across different backbones and categories. Figure 11 illustrates the performance of OD models with different denoising preprocessing under adverse weather. Models without denoising exhibit numerous false positives and missed detections. Compared with 3D‑OutDet, our TripleMixer produces cleaner point clouds, enabling OD models to achieve higher detection accuracy and reducing both false positives and missed detections.
Figure 11: Qualitative results of the OD benchmark under adverse weather. The “Denoising” rows display outputs from different denoising models, and the subsequent rows present object detection results on the corresponding denoised point clouds. Red boxes indicate ground‑truth objects, green boxes denote predicted objects. Turquoise circles highlight incorrect detections, and pink circles indicate missed detections.
Table XI: Ablation Results for Each Layer of TripleMixer.
@@ V-D Object Detection (OD) Benchmark
Table XII: Ablation Results for XY Projection Resolutions with Z Resolution Fixed at 32.
@@
Table XIII: Ablation Results for Z Projection Resolutions with XY Resolutions Fixed at [256, 256]
@@
Table XIV: Ablation Results for Wavelet Decomposition Levels in FMX Layer
@@
VI Discussion
VI-A Ablation Study
This section presents ablation studies to evaluate the impact of key components in the TripleMixer model, including three primary layers, projection plane resolution, wavelet decomposition levels and loss functions. All ablation experiments are conducted on the real-world WADS[18] dataset, with same experimental settings as in Section V.
@@ VI-A Ablation Study
Effects of Each Layer. To testify the effectiveness of each Layer of TripleMixer, we conduct an extensive ablation study and list the result in Table XI. Experiments I-III represent the accuracy of our TripleMixer model after sequentially removing the GMX Layer, the FMX Layer, and the CMX Layer, respectively. From Experiments I and IV, we observe that the GMX Layer contributes significantly, resulting in F1 and mIoU scores of 0.49 and 0.90, respectively. Similarly, Experiments II and IV indicate that the FMX Layer enhances performance, yielding F1 and mIoU improvements of 0.65 and 1.19, respectively. Finally, the impact of the CMX Layer is evident from Experiments III and IV, showing F1 and mIoU increases of 0.12 and 0.25, respectively. These results demonstrate that the GMX Layer, the FMX Layer, and the CMX Layer are all crucial for improving model performance, each contributing significantly to the overall effectiveness of our TripleMixer model.
Effects of Projection Plane Resolution. In the FMX layer, 3D geometric features are projected onto three 2D planes along the X, Y, and Z axes. Adaptive wavelet decomposition is then applied to extract multi-scale features from each plane. The initial resolution of these projections plays a key role in determining the quality of the extracted features. To evaluate its impact, we conduct ablation studies on different projection resolutions. Considering the asymmetric LiDAR scanning ranges, we separately investigate the XY and Z directions. First, we fix the Z-direction resolution at 32 and explore various resolutions in the XY plane. As shown in Table LABEL:plane_reso_xy, the highest mIoU is achieved at [256, 256], while both lower ([64, 64]) and higher ([512, 512]) resolutions lead to performance drops. Based on this, we fix the XY resolution to [256, 256] and examine different Z-direction resolutions. Table LABEL:plane_reso_z shows that a resolution of 32 again yields the best performance, with other values resulting in suboptimal accuracy. These findings confirm that the projection resolution of [256, 256, 32] provides the best balance between feature granularity and computational efficiency, and is adopted in all practical experiments.
Effects of Wavelet Decomposition Levels. We evaluate the impact of different wavelet decomposition levels in the Frequency Mixer (FMX) on model performance. With the projection resolution fixed at [256, 256, 32] along the XYZ axes, we set the maximum wavelet decomposition level to 3, as higher levels reduce the Z-direction resolution to less than or equal to 2, limiting the extraction of meaningful features. We examine decomposition levels ranging from 0 to 3. As shown in Table LABEL:wavelet, omitting wavelet decomposition yields the lowest accuracy, underscoring its importance in feature extraction. While moderate decomposition improves performance, excessive decomposition introduces information loss. The best results are obtained when the wavelet decomposition level is set to 2, confirming its effectiveness in balancing detail preservation and multi-scale representation.
@@
Effects of Loss Functions. Our model’s loss function primarily consists of three components: the cross-entropy loss, the lovasz loss, and the wavelet regularization terms. We use the cross-entropy loss as the baseline to test the impact of different loss functions on model accuracy. As shown in Table XV, after adding only the lovasz loss, the mIoU and F1 scores of our TripleMixer model improved by 0.37 and 0.21, respectively. Additionally, after incorporating only the wavelet regularization terms, the model’s mIoU and F1 scores increased by 0.26 and 0.14, respectively. When using the combination of all three loss functions, our denoising model achieved the highest mIoU and F1 scores.
Table XVI: Comparison of Model Complexity, Runtime, and Peak Memory Usage.
Method Parameters (M) Runtime (ms) Memory (GB)
Cylinder3D [45] 49.52 127 0.92
LSK3DNet[46] 28.84 89 0.78
DSOR [18] 253
DROR [25] 199
4DenoiseNet[21] 0.57 96 2.47
3D-OutDet[14] 0.01 82 0.33
TripleMixer (Ours) 1.02 61 0.21
VI-B Efficiency Comparisons
Table XVI compares the model complexity, runtime, and peak memory usage during inference across different methods. General point cloud segmentation models adapted for denoising, such as Cylinder3D [45] and LSK3DNet [46], have substantially higher parameter counts of 49.52M and 28.84M and memory footprints of 0.92GB and 0.78GB. Traditional statistical denoising methods DSOR [18] and DROR [25] are computationally expensive with long runtimes. 3D‑OutDet[14] has the smallest parameter count of 0.01M but a relatively high runtime due to additional k‑NN preprocessing. 4DenoiseNet [21] shows the highest peak memory usage of 2.47GB as it aggregates spatiotemporal features from multiple frames. In contrast, our TripleMixer achieves a favorable balance with a runtime of 61ms and the lowest peak memory usage of 0.21GB, demonstrating high efficiency.
VII Conclusion
In this work, we proposed TripleMixer, a robust plug‑and‑play denoising network that can be seamlessly integrated into LiDAR perception pipelines. We also introduced two large‑scale simulated LiDAR datasets, Weather‑KITTI and Weather‑NuScenes, encompassing diverse adverse weather conditions. Building on these contributions, we established four benchmarks: Denoising, Semantic Segmentation (SS), Place Recognition (PR), and Object Detection (OD), providing a unified framework for evaluating denoising models and their impact on downstream perception tasks. Our model was trained exclusively on the proposed datasets, and all perception models were directly evaluated on real‑world adverse‑weather datasets without retraining. Extensive experiments demonstrated that TripleMixer achieves state‑of‑the‑art denoising performance and substantially improves SS, PR, and OD accuracy under real‑world adverse weather, underscoring the strong generalization capability of both the proposed model and datasets.
Future Work. While the multi‑branch mixer design of TripleMixer effectively suppresses noise, its relatively high parameter count may hinder deployment in resource‑constrained settings. Future work will focus on developing a more lightweight denoising network. Furthermore, while synthetic weather data provides valuable scalability and controllability, it may not fully capture the subtle complexities of real‑world weather conditions. We will explore more realistic simulation techniques and expand real‑world adverse‑weather datasets to further improve model generalization.